OpenStudy (anonymous):

Please help ! Find dy/dx if y=2x^3/3.

OpenStudy (anonymous):

dy/dx=d/dx(2x^3/3)=(2/3)d/dx(x^3)=(2/3)*3x^2=2x^2

OpenStudy (anonymous):

thanks ! (: